Suppose an indifference curve is given by the equation U=2*C*T. Assume that initially the consumer owns the bundle C = 20, T = 2. What is the Utility value along this indifference curve? Show your calculations! Other than point, (C, T) = (20, 2), list 3 other points on this indifference curve. Graph the indifference curve including the point (C, T) = (20, 2) and the 3 points you listed in part b. Place T on the vertical axis and C on the horizontal axis. Make sure to label the axes, the indifference curve, and the 4 points on the indifference curve.What is the slope of this indifference curve between C=19 and C=20? Show your calculations!
Answer #1

U=2*C*T=2*20*2=80

Equation of IC curve with U=80

40=C*T

Let C=10. ,T=4

C=5. ,T=8

C=4 ,T=10

Slope of indifference between C=19 and C=20 is change of quantity of T to maintain same utility level.

80=2*19*T

T=40/19

∆T=2-40/19=(38-40)/19=-2/19=-0.105

Slope of IC between C=19 and C=20 is -0.105